An improvised billiard ball is used in a local play where a 35 gram billiard ball resting in a corner is suddenly

hit by another 45 grams ball moving 0.5 m/s. After the collision, the 45 grams ball remains at rest while the

other ball moves away. Find the velocity of the ball moving away ?


1
Expert's answer
2022-06-13T08:22:42-0400

Given:

"v_1=0"

"m_1=35\:\rm g"

"v_2=0.5\:\rm m/s"

"m_2=45\:\rm g"

"v_2'=0"


The law of conservation of momentum says

"m_1v_1+m_2v_2=m_1v_1'+m_2v_2'"

Thus

"v_1'=v_1+(v_2-v_2')m_2\/m_1"

"v_1'=0+(0.5-0)45/35=0.64\:\rm m/s"